OpenStudy (anonymous):

find [f*g](x) if f(x)=x^2-5 g(x)=2x

OpenStudy (unklerhaukus):

do you mean \[[f\times g](x)\]or\[[f\circ g](x)\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

2x^2 - 5

OpenStudy (unklerhaukus):

\[f(x)=x^2-5\]\[g(x)=2x\] the composition of \(f\) with \(g\) of \(x\) \[[f\circ g](x)=f\big(g(x)\big)=f(2x)=(2x)^2-5=\]